OpenStudy (anonymous):

HELP with CHAIN RULE!!!!!! y= e^(x^2 +1) * sinx^3

OpenStudy (anonymous):

looks like you need the product rule

OpenStudy (anonymous):

yes but you wud first use chain rule wouldn't you?

OpenStudy (anonymous):

\[\left(fg\right)'=f'g+g'f\] with \[f(x)=e^{2x+1},f'(x)=2e^{2x+1},g(x)=\sin(x^3), g'(x)=3x^2\cos(x^3)\]

OpenStudy (anonymous):

yeah you do need the chain rule to find the derivative of \(e^{2x+}\) and \(\sin(x^3)\)

OpenStudy (anonymous):

oh wow i never thought of it that way! So now use product rule?

OpenStudy (anonymous):

except why wud f(x) be 2x? it wud be x^2 right?

OpenStudy (anonymous):

like it would be e to the power of x^2 +1 not 2x +1

OpenStudy (anonymous):

@satellite73

OpenStudy (amistre64):

consider the chain rule as:\[y(x)=f(u)\] \[\frac{dy}{dx}\frac{dx}{dx}=\frac{df}{du}\frac{du}{dx}\frac{dx}{dx}\]

OpenStudy (amistre64):

for example: \[y = e^{x^2 +1} \] \[\frac{dy}{dx} = \frac{d(e^{x^2+1})}{d(x^2+1)}\frac{d(x^2+1)}{dx}\] if we let u = x^2+1 \[\frac{dy}{dx} = \frac{d(e^{u})}{du}\frac{du}{dx}\]

OpenStudy (amistre64):

y' = 2x e^u